# SmartAgent: Chain-of-User-Thought for Embodied Personalized Agent in Cyber World

### v2

@@ 6 Evaluation
6.1 Embodied Personalized Reasoning
We first investigate the comprehensive capabilities of SmartAgent in handling embodied tasks and personalized reasoning, primarily focusing on SmartSpot for validation. Specifically, we categorize the tasks based on the SmartSpot into simpler single-domain tasks and more complicated scenarios that mimic real multi-channel interactions. We exclude the MEDICINE channel to evaluate the zero-shot ability later. We select SeeClick [3] (the current SOTA GUI agent based on pure visual observation), along with Qwen-VL [1] and Qwen2-VL [38], as our baselines.
Results & analysis. Tab. 2 shows SmartAgent’s average performance across all channels. It achieved the best results in the Ele.Acc metric and ranked second in SSR, indicating strong foundational embodied perception ability. SmartAgent inferred 71% of users’ underlying requirements, afterward leading to a 24% accuracy in item recommendations. More importantly, Fig. 4 highlights that SmartAgent consistently outperformed in all multi-channel scenarios, whereas specialized GUI agents and general LLMs showed slightly weaker embodied outputs. Overall, these evaluations confirm SmartAgent’s comprehensive capabilities on both embodied and personalized tasks. Qwen2-VL had unexpectedly poor results and will not be discussed further.
@@ 6.2 GUI Grounding
An important consideration is whether training with personalized capabilities leads to catastrophic forgetting of pre-trained embodied abilities. In this section, we evaluate SmartAgent on a renowned GUI Grounding benchmark SmartSpot to assess its foundational perception of raw GUI data. The comparative baselines are divided into two main categories: GUI specialist models and general LLMs.
Results & analysis. As shown in Appendix Tab. 5, SmartAgent achieves second-best results in most metrics, even securing the top position in Desktop-Icon/Widget. This indicates that training with personalized capabilities not only preserves foundational embodied abilities but also enhances its proficiency in operations that involve user intent. This is also the primary achievement of the COUT paradigm.
6.3 Autonomous GUI Operation
Automated execution of human instructions is a foundational capability of embodied agents. In this section, we validate SmartAgent’s basic abilities to handle GUI action episodes autonomously using the classic GUI agent benchmark Mind2Web.
Results & analysis. We transfer the well-trained SmartAgent on SmartSpot to Mind2Web. As shown in Appendix Tab. 6, it achieves second place generally in the pure vision-based baselines across three downstream scenarios. This indicates that SmartAgent possesses strong embodied generalization capabilities, enabling seamless integration into various downstream tasks.
@@ 6.4 Explicit & Implicit Personalized Reasoning
In this section, we present a case study to demonstrate SmartAgent’s performance in explicit user underlying reasoning and implicit item recommendation tasks. As shown in Fig. 5, SmartAgent correctly predicted the action of clicking on the economy class option in the visual observation. This prediction is reflected in the summarization of the user’s underlying requirements. Leveraging this textual representation of explicit needs, SmartAgent made recommendations for the last two flight options, effectively addressing the user’s implicit requirements.
6.5 Zero-shot Reasoning
Results & analysis. As shown in Tab. 3, SmartAgent surprisingly exceeds its average performance achieved through fine-tuning on SmartSpot in the Exp.Acc metric. This demonstrates SmartAgent’s robustness in interpreting users’ explicit intentions and also indicates a preliminary zero-shot reasoning capability.
Table 3: Results of zero-shot reasoning on channel MEDICINE.
Type MEDICINE
Ele.Acc SSR Exp.Acc Imp.Acc
Zero-shot 0.04 0.08 0.77 0.14
Full-stage 0.64 0.50 0.71 0.24
6.6 More Insights into COUT
In this section, we provide deeper insights into COUT reasoning, focusing on the two-stage training and end-to-end training settings (without underlying thought generation). As shown in Tab. 4, SmartAgent with end-to-end training exhibits slightly better embodied performance, although it is limited in user underlying perceptions. We therefore argue that balancing embodied perceptions in response to changing environments with reliable personal service is crucial for future COUT research.
Table 4: Results of end-to-end and two-stage training.
Ele.Acc SSR Exp.Acc Imp.Acc
Two-stage 0.64 0.50 0.71 0.24
End-to-End 0.67 0.53 0.66 0.31
7 Conclusion and Future Work
In this paper, we introduce a novel embodied reasoning paradigm, COUT, which for the first time defines an embodied personalized task. We establish a clear definition and components of the COUT paradigm and analyze its challenges. To address these challenges, we propose SmartAgent to instantiate COUT through a two-stage training from essential GUI reasoning to high-level user thought reasoning. For benchmarking this progress in SmartAgent, we created SmartSpot, the first embodied AI benchmark featuring explicit personalization evaluations. Results on SmartSpot demonstrate the effectiveness and proficiency of SmartAgent over full-stage embodied personalized reasoning tasks. Furthermore, SmartAgent showcases the key capability of zero-shot embodied reasoning, highlighting its potential for efficient adaptation in new scenarios.
Supplementary Material
{strip}
@@ 7 Conclusion and Future Work
Table 6: Autonomous GUI operation results on Mind2Web.
Methods Pure Visual Cross-Task Cross-Website Cross-Domain
Ele.Acc Op.F1 SSR Ele.Acc Op.F1 SSR Ele.Acc Op.F1 SSR
MindAct (gen) 20.2 52.0 17.5 13.9 44.7 11.0 14.2 44.7 11.9
MindAct 5.1 75.7 52.0 42.0 65.2 38.9 42.1 66.5 39.6
GPT-3.5-Turbo 20.3 56.6 17.4 19.3 48.8 16.2 21.6 52.8 18.6
GPT-4 41.6 60.6 36.2 35.8 51.1 30.1 37.1 46.5 26.4
Qwen-VL 15.9 86.7 13.3 13.2 83.5 9.2 14.1 84.3 12.0
SeeClick 28.3 87.0 25.5 21.4 80.6 16.4 23.2 84.8 20.8
SmartAgent 24.5 80.2 20.6 18.9 74.9 15.0 20.3 77.9 17.1
8 Action Space of SmartSpot
SmartSpot offers common GUI actions in daily usage. We follow AiTW [31] to assign a action_type_id to each action for model prediction. The full action space is shown as follows:
  • type("text"):3. An action that types text at (x,y).
  • scroll(direction): Scroll actions for the screen, scroll up/down/left/right are assigned with 1, 0, 8, and 9.
  • complete:10. An action of determining if has reached the item pool.
  • BACK:5. An action for returning to the previous step.
  • HOME:6. An action for returning to the homepage.
9 More Training Details
Following the approach of [3], We intuitively present numerical coordinates as natural languages without additional tokenization or pre-/post-processing. We train SmartAgent 15 epochs for both the embodied and personalized stages. Embodied baselines are trained for 15 rounds. Results on the ScreenSpot and Mind2Web benchmarks are evaluated via direct testing and 10 epochs of training, as in [3].
10 More Results
This section shows SmartAgent’s performance results on basic embodied sub-tasks. The results of GUI grounding on the benchmark ScreenSpot are presented in Tab. 5, while the results of autonomous GUI operation on the benchmark Mind2Web can be found in Tab. 6.

### v3

@@ 6.2 GUI Grounding
An important consideration is whether training with personalized capabilities leads to catastrophic forgetting of pre-trained embodied abilities. In this section, we evaluate SmartAgent on a renowned GUI Grounding benchmark ScreenSpot [3] to assess its foundational perception of raw GUI data. The comparative baselines are divided into two main categories: GUI specialist models and general LLMs.
Results & analysis. As shown in Table 3, SmartAgent achieves second-best results in most metrics, even securing the top position in Desktop-Icon/Widget. This indicates that training with personalized capabilities not only preserves foundational embodied abilities but also enhances proficiency in operations involving user intent. This is also the primary achievement of the COUT paradigm. Note that we do not overly require SmartAgent to maintain a SOTA level on such traditional test environment that do not include personalized factors. Our primary goal in this section is to ensure that it retains sufficient basic capabilities in tasks after incorporating personalized considerations.
6.3 Autonomous GUI Operation
@@
Automated execution of human instructions is a foundational capability of embodied agents. In this section, we validate SmartAgent’s basic abilities to handle GUI action episodes autonomously using the classic GUI agent benchmark Mind2Web [6].
Results & analysis. We transfer the well-trained SmartAgent on SmartSpot to Mind2Web. As shown in Appendix Table 6, it achieves second place generally in the pure vision-based baselines across three downstream scenarios. This indicates that SmartAgent possesses strong embodied generalization capabilities, enabling seamless integration into various downstream tasks. Similarly, we do not expect SmartAgent to surpass the SOTA model in this basic task. The first-tier autonomous agent capabilities that have not been forgotten in the backbone model have laid the foundation for subsequent personalized reasoning tasks.
@@ 6.4 Explicit & Implicit Personalized Reasoning
In this section, we present a case study to demonstrate SmartAgent’s performance in explicit user underlying reasoning and implicit item recommendation tasks. As shown in Figure 5, SmartAgent correctly predicted the action of clicking on the economy class option in the visual observation. This prediction is reflected in the summarization of the user’s underlying requirements. Leveraging this textual representation of explicit needs, SmartAgent made recommendations for the last two flight options, effectively addressing the user’s implicit requirements.
@@
Table 4: Results of zero-shot reasoning on channel MEDICINE.
Type MEDICINE
Ele.Acc SSR Exp.Acc Imp.Acc
Zero-shot 0.04 0.08 0.77 0.14
Full-stage 0.64 0.50 0.71 0.24
Table 5: Results of end-to-end training and two-stage training.
Ele.Acc SSR Exp.Acc Imp.Acc
Two-stage 0.64 0.50 0.71 0.24
End-to-End 0.67 0.53 0.66 0.31
6.5 Zero-shot Reasoning
Results & analysis. As shown in Table 4, SmartAgent surprisingly exceeds its average performance achieved through fine-tuning on SmartSpot in the Exp.Acc metric. We argue that in cold-start scenarios like MEDICINE, the training data may be very limited, rendering fine-tuning ineffective or even leading to overfitting. SmartAgent’s zero-shot capability can compensate for this limitation by leveraging general GUI knowledge to interact directly with new scenarios. This enhancement is also evident in the subsequent item recommendation task, which approaches the performance of full-stage fine-tuning. Overall, this demonstrates SmartAgent’s robustness in interpreting users’ explicit intentions and also indicates a preliminary zero-shot reasoning capability.
6.6 More Insights into COUT
In this section, we provide deeper insights into COUT reasoning, focusing on the two-stage training and end-to-end training settings (without underlying thought generation). As shown in Table 5, SmartAgent with end-to-end training exhibits even slightly better embodied performance, although it is limited in user underlying perceptions. This may be due to the large number of long action episodes, which forces the model to overfit underlying requirements or even accumulate serious errors, which then negatively affects the subsequent personalized stages. We, therefore, argue that balancing embodied perceptions in response to changing environments with reliable personal service is crucial for future COUT research.
7 Conclusion and Future Work
In this paper, we introduce a novel embodied reasoning paradigm, COUT, which for the first time defines an embodied personalized task. We establish a clear definition and components of the COUT paradigm and analyze its challenges. To address these challenges, we propose SmartAgent to instantiate COUT through a two-stage training from essential GUI reasoning to high-level user thought reasoning. To benchmark this progress in SmartAgent, we created SmartSpot, the first embodied AI benchmark featuring explicit personalization evaluations. Results on SmartSpot demonstrate the effectiveness and proficiency of SmartAgent over full-stage embodied personalized reasoning tasks. Furthermore, SmartAgent showcases the key capability of zero-shot embodied reasoning, highlighting its potential for efficient adaptation in new scenarios.
As for the future work, we plan to further extend the scale of the benchmark, with more extensive experiments and analysis. We also plan to apply the proposed method to the real user devices.